Stephanie bought 3 textbooks for her Chemistry, math and Ceramics courses. The chemistry book cost 20.00 dollars more than the ceramics book. the Math book cost twice as much as the chemistry book. If the total bill was 180.00 what is the price of each book?
Answer by riyotaku(1)

Let ch=chem, m=math, c=ceramics
therefore, ch=c+20 and m=2(c+20)
and thus ch+m+c= 2(c+20) + c+20 + c
180= 3(c+20) + c
180= 3c + 60 + c
180= 4c + 60
180 - 60= 4c
60= 4c
60/4= c
and therefore c= $15
since we have found c we can then find ch and m:
ch=c + 20
=15 + 20
=$35
and
m= 2(c + 20)
= 2(15 + 20)
= 2 x 35
= $70